# Searches for the BSM scenarios at the LHC using decision tree based machine learning algorithms: A comparative study and review of Random Forest, Adaboost, XGboost and LightGBM frameworks

### v1

@@ 1 Introduction
It is important to note that the accumulated luminosity of Run-II data is approximately $\sim$ 140 $\rm{fb}^{-1}$, which is about 5% of the luminosity of planned upgrade of high luminosity LHC (HL-LHC) run ($\mathcal{L}=3000\rm{fb}^{-1}$). The LHC Collaboration initiated the Run-III operation in July 2022, and it will continue the operation until the planned three-year-long shutdown (LS3) in preparation for the high luminosity upgrade, which is scheduled to commence in 2026. The constraints on sparticle masses, derived from LHC Run-I and Run-II data, are primarily derived within the framework of simplified SUSY scenarios, which involve specific assumptions regarding decay modes, branching ratios, etc. Also, the majority of the analyses have concentrated on prompt decay scenarios. However, there are several pockets of SUSY parameter spaces with light sparticles that remain less explored, e.g., quasi-degenerate SUSY scenarios where the next to LSP (NLSP) - LSP mass gap is very small Bhattacherjee:2013wna ; Dutta:2015exw ; Chakraborti:2017dpu ; Chowdhury:2016qnz ; Dutta:2017jpe . Such compressed SUSY models are also highly motivated in the context of Dark Matter (DM) relic density KumarBarman:2020ylm ; Barman:2022jdg ; He:2023lgi ; Chakraborti:2015mra ; Chakraborti:2014gea ; Chakraborti:2017dpu ; Chowdhury:2016qnz ; Bhattacharyya:2011se ; Choudhury:2012tc ; Baer:2021aax , Muon (g-2) anomaly Chakraborti:2022vds ; He:2023lgi ; Chakraborti:2015mra ; Chakraborti:2014gea ; Baer:2021aax ; Athron:2021iuf ; Endo:2021zal ; Chakraborti:2021bmv ; Choudhury:2017acn ; Choudhury:2017fuu ; Banerjee:2018eaf ; Banerjee:2020zvi ; Chakraborti:2021dli ; Frank:2021nkq ; Ali:2021kxa ; Kowalska:2015zja ; Chakrabortty:2015ika ; Choudhury:2016lku . It has been observed that conventional cut and count analyses are highly competitive for non-degenerate SUSY searches compared to machine learning (ML) analyses. But ML based analyses give better sensitivity in both degenerate and non-degenerate BSM physics searches.
@@
The term “machine learning” (ML) was first proposed by A. Samuel in 1959 Samuel1988 . After six decades, ML algorithms are now being widely used in our daily lives and across various industries, e.g., filtering email, social media recommendation, cyber security, image analysis and disease detection in healthcare, data analysis in the finance sector, autonomous vehicles, marketing and advertising, etc. In the future, the applications of ML will continue to expand rapidly with the advancement of technology and the availability of more data. high energy physics (HEP) experimental and phenomenological analyses deal with large amount of data and it has been observed that the use of ML techniques leads to improvements in the data analyses of several HEP fields. For particle identification, event selection, object reconstruction, event classification, etc., the experimental high energy physics (HEP-Ex) collaborations have been using the conventional ML algorithms for more than three decades. boosted decision tree (BDT) is one of the most popular algorithms which has been widely used by the HEP-Ex collaborations e.g., the search of single top quark prediction by the CDF and D0 collaborations CDF:2010eor ; D0:2008wma ; D0:2006ngk , Higgs discovery at the LHC CMS:2014afl . For a long time, the HEP community has used BDT and other algorithms implemented in the Toolkit for Multivariate Data Analysis (TMVA) software package TMVA:2007ngy while XGBoost (Extreme Gradient Boosting) has gained immense popularity in recent years Chen:2016btl . However, it should be noted that the deep neural networks (DNN) or deep learning (DL) techniques, which are based on multilayer NN, are becoming more popular nowadays doi:10.1142/S0217751X20020030 ; dnn ; Bhat:2010zz ; Guest:2018yhq ; Bourilkov:2019yoi ; Schwartz:2021ftp ; Carleo:2019ptp ; Shlomi:2020gdn ; Abdughani:2019wuv . There are several broad reviews in the literature on the applications of BDT Bhat:2010zz ; Cornell:2021gut ; Coadou:2022nsh , deep learning algorithm doi:10.1142/S0217751X20020030 ; dnn ; Bhat:2010zz ; Guest:2018yhq ; Bourilkov:2019yoi ; Schwartz:2021ftp ; Carleo:2019ptp ; Shlomi:2020gdn ; Abdughani:2019wuv like convolutional neural networks (CNN), recurrent neural networks (RNN) etc. in the context of high energy physics.
@@
The collider analyses in HEP typically involve with searches of new physics signals and the precise measurements of the existing known SM processes. To achieve this, one has to look for faint signal from a large amount of background where the distribution of signal and backgrounds have significant overlap. Traditional cut-based analysis shows less sensitivity for significant overlap scenarios. In such cases, the ML algorithms are more powerful in discriminating signals from backgrounds. In this article, we will focus on the various kinds of decision tree (DT) based algorithms and study how these algorithms lead to improvement in the search sensitivity compared to cut-based methods considering a case study of electroweakino pair production at the HL-LHC. The structure of the article is as follows. In Sec. 2, we introduce the basic concepts of machine learning along with different kinds of metrics. In Sec. 3, we briefly summarize the major important analyses involving the use of ML algorithms with an emphasis on DT-based algorithms. The basic concepts of decision trees are discussed in Sec. 4. Also, we present a concise overview of Random Forest, AdaBoost, and extreme gradient boosting algorithms such as XGBoost and LightGBM in this section. In Sec. 5, we investigate the improvement of search sensitivity by these four ML algorithms with comparison to cut-and-count analysis using a SUSY scenario with $\widetilde{\chi}_{1}^{\pm}\widetilde{\chi}_{2}^{0}$ (wino-like) pair production. We also present the role of hyperparameters in different ML framework along with the feature importance study using SHapley values. Finally, we summarize the paper in Sec. 6.
@@ 2 Basic concepts of machine learning
Machine learning, as a subset of artificial intelligence (AI), involves the development of various models that can learn from diverse datasets and execute tasks without requiring explicit programming. To perform a machine learning (ML) analysis, it is essential to collect data, which comprises information about the desired output that we want the computer to learn. The ML algorithms can be broadly categorized into supervised and unsupervised learning inbook ; hastie2009elements ; carbonell1983overview . In addition to these two, there is also a third category known as semi-supervised learning. In supervised learning, each input data point ($x_{i}$, i = 1,….n, for n input data samples) is accompanied by a target variable or output label ($y_{i}$) and the algorithm learns to extrapolate patterns from the provided training data in order to predict the output labels for unseen (testing) data points. In other words, the aim of the supervised learning is to acquire a mapping function $f$ from the input data to output label: $f:x_{i}\rightarrow y_{i}$ such that it can accurately predict $y_{i}$ for a new data sets (testing) which do not have output labels.
@@ 2.1 Loss Function
For binary classification tasks the commonly used the most commonly used function is Binary Cross-Entropy Loss or Log Loss, which is defined as:
@@
where $y_{i}$ is the predicted probability and $\hat{y_{i}}$ is the true binary label (0 or 1). Categorical cross-entropy loss ($\sum_{i=1}^{n}\hat{y_{i}}$ log(y)) is used in multi-class classification tasks.
@@ 2.3 Measures of classification performance
After training, tuning of hyperparameter and testing the HEP data, our objective is to evaluate the effectiveness and accuracy of the machine learning models in solving a specific task, e.g., classify or distinguish between signal and background events despite their similar signature. Performance metrics are quantitative indicators used to measure the model’s ability to make accurate predictions or classifications. By evaluating performance metrics, we can compare different ML models and optimize the hyperparameter tuning to improve the model’s effectiveness. Sometimes, the highly skewed distribution of classes in the dataset leads to class imbalance, which affects the performance of the ML algorithm. Techniques like oversampling of the minority class or undersampling of the majority class, adjustment of class weights, cost-effective learning, etc., can be employed to address the class imbalance in ML Cornell:2021gut ; weiss2003learning ; Chawla2005 ; sulaiman ; satio ; LUQUE2019216 ; 10.1007/978-3-642-04962-0_53 . To evaluate the performance metrics, the algorithm checks the actual and predicted values of observations in the dataset by forming confusion matrix [^3] It is also known as error matrix or classification table.. When the signal/background events are correctly classified as signal/background events, then those events are called True Positive (TP) or True Negative (TN). On the other hand, False Positive (FP), False Negative (FN) are the number of events where the actual background/signal events are incorrectly classified as signal/background events. In the Table. 1, we present the confusion matrix graphically for binary classification. With these definitions, we can define the following various performance metrics to evaluate the performance of the model:
@@
  • Approximate Median Significance (ams) score: In the field of high energy physics, the primary goal is to optimize the discovery significance. To estimate the discovery significance, the formulas $s/\sqrt{b}$ or $s/\sqrt{(s+b)}$ are commonly used [^5] For discovery of a new particle the significance should be $\geq 5\sigma$ and for exclusion it should be $\geq 2\sigma$ , where s and b denote the numbers of signal and background events, respectively, that remain after the signal selection cuts. It may be noted that in a typical Poisson counting experiment, where $n$ events are observed, the Poisson distribution often features a large mean value, $(s+b)$. However, the formula $s/\sqrt{b}$ is valid only for $b>>s$ and it overestimates the discovery significance when the background events are small Cowan2012DiscoverySF ; Cowan:2010js . For a Poisson counting experiment with negligible uncertainty, the Asimov approximation for the median significance (ams score) is given by Cowan2012DiscoverySF ; Cowan:2010js :
@@ 3 Machine learning in High Energy Physics
Jet tagging: In collider experiments, the reconstruction and identification of hadronic jets are integral components for physics analyses. Jets originating from massive particles (e.g., bottom quark, top quark, $W/Z/h$ bosons or supersymmetric partner of top quark ($m_{\widetilde{t_{1}}}$)) are usually boosted and collinear. Jets emerging from b-quark are associated with long lifetime and secondary vertex and can be differentiated from other jets coming from light quarks. Motivated by computer vision techniques, jet images was first introduced in Ref. Cogan:2014oua and then many groups have used modern deep neural network (DNN) architectures for jet tagging with jet images deOliveira:2015xxd ; Baldi:2016fql ; Komiske:2016rsd ; Chakraborty:2019imr ; ATLAS:2017dfg . The CMS collaboration has developed DeepCSV CMS:2017wtu ; Bols:2020bkb , DeepFlavour  2020HeavyFI ; deepflv and DeepJet Bols:2020bkb taggers for multiclass classification for light jets, gluon jets, c-jets and b-jets using a deep learning algorithm. In a recent work ATLAS:2019bwq , the ATLAS Collaboration has developed two high-level b-tagging algorithms - MV2 and DL1 Krizhevsky2012ImageNetCW . MV2 Krizhevsky2012ImageNetCW algorithm is based on BDT  [^7] This same algorithm has been used in another analysis ATLAS:2020hpj , where the ATLAS collaboration has reported the first evidence of $t\bar{t}t\bar{t}$ production. and is trained within the TMVA framework TMVA:2007ngy . The performance of b-jets tagging using the simulated $t\bar{t}$ events has been presented in Fig.1 in ATLAS:2019bwq for BDT and deep feed-forward neural network algorithms. Although for jet images, tagging, and substructure studies, deep learning algorithms are the most efficient ones, the ATLAS collaboration recently showed that W-boson tagging and top quark tagging ML algorithms lead to a significant gain in efficiency compared to cut-based analysis, as presented in Fig.4 ATLAS:2018wis . This analysis has been performed using LHC Run-II data with $\mathcal{L}=36\rm{fb}^{-1}$. It may be noted that both the BDT and DNN-based algorithms perform similarly to each other for all signal efficiencies ATLAS:2018wis .
@@ 3.1.1 Searches for RPC SUSY scenarios using BDT
In the Ref Cornell:2021gut , the authors have shown that gradient boosting algorithm can extend the 95% C.L. exclusion limits derived from traditional cut-based methods using an example of smuon pair production at the 13 TeV LHC. The authors have also studied the utility of various performance metrics (auc, ams, F-score etc.) using XGBoost toolkit, feature importance using SHAP package Shapley+1953+307+318 ; lundberg2018consistent in great details Cornell:2021gut . The possibility of constructing a general machine learning model that may be applied to probe a two-dimensional mass plane has also been examined in this work. In Alvestad:2021sje , the authors pointed out that XGBoost algorithm can significantly increase the detectability of a SUSY models with a gravitino type LSP and a metastable sneutrino NLSP by analyzing the events coming from the electroweakinos production along with the slepton productions. The possibility of probing the light Higgsinos, which are still allowed by LHC Run-II and LZ experiment data, at the upcoming HL-LHC run via XGBoost framework has been studied in Ref. Barman:2024xlc

### v2

@@ 1 Introduction
It is important to note that the accumulated luminosity of Run-II data is approximately $\sim$ 140 $\rm{fb}^{-1}$, which is about 5% of the luminosity of planned upgrade of high luminosity LHC (HL-LHC) run ($\mathcal{L}=3000\rm{fb}^{-1}$). The LHC Collaboration initiated the Run-III operation in July 2022, and it will continue the operation until the planned three-year-long shutdown (LS3) in preparation for the high luminosity upgrade, which is scheduled to commence in 2026. The constraints on sparticle masses, derived from LHC Run-I and Run-II data, are primarily derived within the framework of simplified SUSY scenarios, which involve specific assumptions regarding decay modes, branching ratios, etc. Also, the majority of the analyses have concentrated on prompt decay scenarios. However, there are several pockets of SUSY parameter spaces with light sparticles that remain less explored, e.g., quasi-degenerate SUSY scenarios where the next to LSP (NLSP) - LSP mass gap is very small Bhattacherjee:2013wna ; Dutta:2015exw ; Chakraborti:2017dpu ; Chowdhury:2016qnz ; Dutta:2017jpe . Such compressed SUSY models are also highly motivated in the context of Dark Matter (DM) relic density KumarBarman:2020ylm ; Barman:2022jdg ; He:2023lgi ; Chakraborti:2015mra ; Chakraborti:2014gea ; Chakraborti:2017dpu ; Chowdhury:2016qnz ; Bhattacharyya:2011se ; Choudhury:2012tc ; Baer:2021aax , Muon (g-2) anomaly Chakraborti:2022vds ; He:2023lgi ; Chakraborti:2015mra ; Chakraborti:2014gea ; Baer:2021aax ; Athron:2021iuf ; Endo:2021zal ; Chakraborti:2021bmv ; Choudhury:2017acn ; Choudhury:2017fuu ; Banerjee:2018eaf ; Banerjee:2020zvi ; Chakraborti:2021dli ; Frank:2021nkq ; Ali:2021kxa ; Kowalska:2015zja ; Chakrabortty:2015ika ; Choudhury:2016lku ; Cao:2022htd ; Cao:2023juc . It has been observed that conventional cut and count analyses are highly competitive for non-degenerate SUSY searches compared to machine learning (ML) analyses. However, ML-based analyses have better sensitivity in both degenerate and non-degenerate BSM physics searches.
@@
The term “machine learning” (ML) was first proposed by A. Samuel in 1959 Samuel1988 . After six decades, ML algorithms are now being widely used in our daily lives and across various industries, e.g., filtering email, social media recommendation, cyber security, image analysis and disease detection in healthcare, data analysis in the finance sector, autonomous vehicles, marketing and advertising, etc. In the future, the applications of ML will continue to expand rapidly with the advancement of technology and the availability of more data. high energy physics (HEP) experimental and phenomenological analyses deal with large amount of data and it has been observed that the use of ML techniques leads to improvements in the data analyses of several HEP fields. For particle identification, event selection, object reconstruction, event classification, etc., the experimental high energy physics (HEP-Ex) collaborations have been using the conventional ML algorithms for more than three decades. Boosted decision tree (BDT) is one of the most popular algorithms which has been widely used by the HEP-Ex collaborations, e.g., the search of single top quark prediction by the CDF and D0 collaborations CDF:2010eor ; D0:2008wma ; D0:2006ngk , Higgs discovery at the LHC CMS:2014afl . For a long time, the HEP community has used BDT and other algorithms implemented in the Toolkit for Multivariate Data Analysis (TMVA) software package TMVA:2007ngy while XGBoost (Extreme Gradient Boosting) has gained immense popularity in recent years Chen:2016btl . However, it should be noted that the deep neural networks (DNN) or deep learning (DL) techniques, which are based on multilayer NN, are becoming more popular nowadays doi:10.1142/S0217751X20020030 ; dnn ; Bhat:2010zz ; Guest:2018yhq ; Bourilkov:2019yoi ; Schwartz:2021ftp ; Carleo:2019ptp ; Shlomi:2020gdn ; Abdughani:2019wuv ; Hammad:2023sbd ; Hammad:2024cae ; Arganda:2024eub . There are several broad reviews in the literature on the applications of BDT Bhat:2010zz ; Cornell:2021gut ; Coadou:2022nsh ; Arganda:2024eub , deep learning algorithm doi:10.1142/S0217751X20020030 ; dnn ; Bhat:2010zz ; Guest:2018yhq ; Bourilkov:2019yoi ; Schwartz:2021ftp ; Carleo:2019ptp ; Shlomi:2020gdn ; Abdughani:2019wuv ; Hammad:2023sbd ; Hammad:2024cae ; Arganda:2024eub like convolutional neural networks (CNN), recurrent neural networks (RNN) etc. in the context of high energy physics.
@@
The collider analyses in HEP typically involve searches of new physics signals and the precise measurements of the existing known SM processes. To achieve this, one has to look for faint signals from a large amount of background where the distribution of signal and backgrounds have significant overlap. Traditional cut-based analysis shows less sensitivity for significant overlap scenarios. In such cases, the ML algorithms are more powerful in discriminating signals from backgrounds. In this article, we will focus on the various kinds of decision tree (DT) based algorithms and study how these algorithms lead to improvement in the search sensitivity compared to cut-based methods considering a case study of electroweakino pair production at the HL-LHC. The structure of the article is as follows. In Sec. 2, we introduce the basic concepts of machine learning along with different kinds of metrics. In Sec. 3, we briefly summarize the major important analyses involving the use of ML algorithms with an emphasis on DT-based algorithms. The basic concepts of decision trees are discussed in Sec. 4. Also, we present a concise overview of Random Forest, AdaBoost, and extreme gradient boosting algorithms such as XGBoost and LightGBM in this section. In Sec. 5, we investigate the improvement of search sensitivity by these four ML algorithms with comparison to cut-and-count analysis using a SUSY scenario with $\widetilde{\chi}_{1}^{\pm}\widetilde{\chi}_{2}^{0}$ (wino-like) pair production. We also present the role of hyperparameters in different ML frameworks along with the feature importance study using Shapley values. Finally, we summarize the paper in Sec. 6.
@@ 2 Basic concepts of machine learning
Machine learning, as a subspace of artificial intelligence (AI), involves the cultivation of various models that can learn from diverse datasets and execute tasks without requiring explicit programming. To perform a machine learning (ML) analysis, it is essential to collect data, which comprises information about the desired output that we want the computer to learn. The ML algorithms can be broadly categorized into supervised and unsupervised learning inbook ; hastie2009elements ; carbonell1983overview . In addition to these two, there is also a third category known as semi-supervised learning. In supervised learning, each input data point ($x_{i}$, i = 1,….n, for n input data samples) is accompanied by a target variable or output label ($y_{i}$) and the algorithm learns to extrapolate patterns from the provided training data in order to predict the output labels for unseen (testing) data points. In other words, the aim of the supervised learning is to acquire a mapping function $f$ from the input data to output label: $f:x_{i}\rightarrow y_{i}$ such that it can accurately predict $y_{i}$ for a new data sets (testing) which do not have output labels.
@@ 2.1 Loss Function
For binary classification tasks the most commonly used loss function is Binary Cross-Entropy Loss or Log Loss, which is defined as:
@@
where $y_{i}$ is the predicted probability and $\hat{y_{i}}$ is the true binary label (0 or 1). Categorical cross-entropy loss ($\sum_{i=1}^{n}\hat{y_{i}}$ log(y)) is used in multi-class classification tasks. The gradient descent technique is used to find the local minima of loss function by an iterative approach. By the iterative method it finds the optimized parameters and improves the performance of the model. Let us consider a loss function $L$, which is differentiable in a neighborhood of a point $p$. $L(p)$ will decrease the fastest in the direction of the loss function ($-\nabla L(p)$). For two consecutive steps, say $n$ and $n+1$, one can write,
$p_{n+1}=p_{n}-\eta\nabla L(p_{n})$ (4)
where $\eta$ is the learning rate. It is worth noting that the choice of $\eta$ is crucial for function optimization. If $\eta$ is too big, it may not reach the minimum and can bounce back and forth.
@@ 2.3 Measures of classification performance
After training, tuning the hyperparameter and testing the HEP data, our objective is to evaluate the effectiveness and accuracy of the machine learning models in solving a specific task, e.g., classify or distinguish between signal and background events despite their similar signature. Performance metrics are quantitative indicators used to measure the model’s ability to make accurate predictions or classifications. By evaluating performance metrics, we can compare different ML models and optimize the hyperparameter tuning to improve the model’s effectiveness. Sometimes, the highly skewed distribution of classes in the dataset leads to class imbalance, which affects the performance of the ML algorithm. Techniques like oversampling of the minority class or undersampling of the majority class, adjustment of class weights, cost-effective learning, etc., can be employed to address the class imbalance in ML Cornell:2021gut ; weiss2003learning ; Chawla2005 ; sulaiman ; satio ; LUQUE2019216 ; 10.1007/978-3-642-04962-0_53 . To evaluate the performance metrics, the algorithm checks the actual and predicted values of observations in the dataset by forming confusion matrix [^3] It is also known as error matrix or classification table.. When the signal/background events are correctly classified as signal/background events, then those events are called True Positive (TP) or True Negative (TN). On the other hand, False Positive (FP), False Negative (FN) are the number of events where the actual background/signal events are incorrectly classified as signal/background events. In the Table. 1, we present the confusion matrix graphically for binary classification. With these definitions, we can define the following various performance metrics to evaluate the performance of the model:
@@
  • Approximate Median Significance (ams) score: In the context high energy physics, the primary objective is to optimize the discovery significance. To estimate the discovery significance, the formulas $s/\sqrt{b}$ or $s/\sqrt{(s+b)}$ are commonly used [^5] For discovery of a new particle the significance should be $\geq 5\sigma$ and for exclusion it should be $\geq 2\sigma$ , where s and b denote the numbers of signal and background events, respectively, that remain after the signal selection cuts. It may be noted that in a typical Poisson counting experiment, where $n$ events are observed, the Poisson distribution often features a large mean value, $(s+b)$. However, the formula $s/\sqrt{b}$ is valid only for $b>>s$ and it overestimates the discovery significance when the background events are small Cowan2012DiscoverySF ; Cowan:2010js . For a Poisson counting experiment with negligible uncertainty, the Asimov approximation for the median significance (ams score) is given by Cowan2012DiscoverySF ; Cowan:2010js :
@@ 3 Machine learning in High Energy Physics
Jet tagging: In collider experiments, the reconstruction and identification of hadronic jets are integral components of physics analyses. Jets originating from massive particles (e.g., bottom quark, top quark, $W/Z/h$ bosons or supersymmetric partner of the top quark ($m_{\widetilde{t_{1}}}$)) are usually boosted and collinear. Jets emerging from b-quark are associated with long lifetime and secondary vertex and can be differentiated from other jets coming from light quarks. Motivated by computer vision techniques, jet images was first introduced in Ref. Cogan:2014oua and then many groups have used modern deep neural network (DNN) architectures for jet tagging with jet images deOliveira:2015xxd ; Baldi:2016fql ; Komiske:2016rsd ; Chakraborty:2019imr ; ATLAS:2017dfg . The CMS collaboration has developed DeepCSV CMS:2017wtu ; Bols:2020bkb , DeepFlavour  2020HeavyFI ; deepflv and DeepJet Bols:2020bkb taggers for multiclass classification for light jets, gluon jets, c-jets and b-jets using a deep learning algorithm. In a recent work ATLAS:2019bwq , the ATLAS Collaboration has developed two high-level b-tagging algorithms - MV2 and DL1 Krizhevsky2012ImageNetCW . MV2 Krizhevsky2012ImageNetCW algorithm is based on BDT  [^7] This same algorithm has been used in another analysis ATLAS:2020hpj , where the ATLAS collaboration has reported the first evidence of $t\bar{t}t\bar{t}$ production. and is trained within the TMVA framework TMVA:2007ngy . The performance of b-jets tagging using the simulated $t\bar{t}$ events has been presented in Fig.1 in ATLAS:2019bwq for BDT and deep feed-forward neural network algorithms. Although for jet images, tagging, and substructure studies, deep learning algorithms are the most efficient ones, the ATLAS collaboration recently showed that W-boson tagging and top quark tagging ML algorithms lead to a significant gain in efficiency compared to cut-based analysis, as presented in Fig.4 ATLAS:2018wis . This analysis has been performed using LHC Run-II data with $\mathcal{L}=36\rm{fb}^{-1}$. It may be noted that both the BDT and DNN-based algorithms perform similarly to each other for all signal efficiencies ATLAS:2018wis .
@@ 3.1.1 Searches for RPC SUSY scenarios using BDT
In the Ref Cornell:2021gut , the authors have shown that the gradient boosting algorithm can extend the 95% C.L. exclusion limits derived from traditional cut-based methods using an example of smuon pair production at the 13 TeV LHC. The authors have also studied the utility of various performance metrics (auc, ams, F-score etc.) using XGBoost toolkit, feature importance using SHAP package Shapley+1953+307+318 ; lundberg2018consistent in great details Cornell:2021gut . The possibility of constructing a general machine learning model that may be applied to probe a two-dimensional mass plane has also been examined in this work. In Alvestad:2021sje , the authors pointed out that XGBoost algorithm can significantly increase the detectability of a SUSY models with a gravitino type LSP and a metastable sneutrino NLSP by analyzing the events coming from the electroweakinos production along with the slepton productions. The possibility of probing the light Higgsinos, which are still allowed by LHC Run-II and LZ experiment data, at the upcoming HL-LHC run via XGBoost framework has been studied in Ref. Barman:2024xlc